Y + 1/2 = - 1/3 (x + 1/2)
I am not understanding solving for y using fractions.

Answer by edjones(8007)

y+ 1/2=-1/3(x+ 1/2)
y+ 1/2=-x/3 - 1/6
6y+3=-2x-1 multiply each side by 6 (LCM).
6y=-2x-4
y=-x/3-2/3
y=-(x+2)/3
.
Ed